cost of a borehole is £400 for the first meter and increases by £25 for each successive meter, calculate the cost of the borehole if it is 24 m deep.
Answer by TimothyLamb(4379)

the problem statement is not clear, so there are two possible interpretations:
---
A) first interpretation:
meter 1 = 400
meter 2 = 25
meter 3 = 25
...
400 + 25*23 = 975
answer A:
borehole cost = 975
---
B) second interpretation:
meter 1 = 400
meter 2 = 400+25
meter 3 = 400+25+25
...
400 + 400+(25*1) + 400+(25*2) ... + 400+(25*22) + 400+(25*23) = ???
24*400 + 25*(1+2+ ... +22+23) = ???
---
the sum of consecutive integers from n to m = (m-n+1)*(n+m)/2
---
24*400 + 25*(m-n+1)*(n+m)/2 = ???
24*400 + 25*(23-1+1)*(1+23)/2 = 16500
answer B:
borehole cost = 16500
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---